### v2

# Fine-Tuning Masked Diffusion for Provable Self-Correction

Abstract
A natural desideratum for generative models is self-correction–detecting and revising low-quality tokens at inference. While Masked Diffusion Models (MDMs) have emerged as a promising approach for generative modeling in discrete spaces, their capacity for self-correction remains poorly understood. Prior attempts to incorporate self-correction into MDMs either require overhauling MDM architectures/training or rely on imprecise proxies for token quality, limiting their applicability. Motivated by this, we introduce PRISM–Plug-in Remasking for Inference-time Self-correction of Masked Diffusions–a lightweight, model-agnostic approach that applies to any pretrained MDM. Theoretically, PRISM defines a self-correction loss that provably learns per-token quality scores, without RL or a verifier. These quality scores are computed in the same forward pass with MDM and used to detect low-quality tokens. Empirically, PRISM advances MDM inference across domains and scales: Sudoku; unconditional text (170M); and code with LLaDA (8B).
1 Introduction
Masked Diffusion Models (MDMs) (sahoo2024simple; gat2024discrete; shi2024simplified) have emerged as a promising approach for generative modeling in discrete domains. At inference time, they start from a fully masked sequence and gradually unmask the masked tokens in arbitrary order to obtain a clean sequence. This flexibility in generation order, combined with recent approaches to deployment at scale (nie2025large; dream2025; gemini2025diffusion; labs2025mercury; gong2025diffucoder), has enabled them to surpass their autoregressive counterparts on several downstream tasks, such as reasoning, coding, and planning.
However, the standard MDM design lacks the ability to self-correct–a natural desideratum for generative models–identifying low-quality or incorrect tokens and correcting them at inference. In MDMs, once a mask token is unmasked, it remains fixed for the entire inference (Figure 1, left). Consequently, the model cannot revise early mistakes, a limitation several recent methods have tried to overcome (campbell2022continuous; lezama2022discrete; zhao2024informed; wang2025remasking; von2025generalized; peng2025pathplanningmaskeddiffusion). Unfortunately, existing approaches come with one of two drawbacks: they either (1) indirectly or inaccurately learn a notion of token quality that lacks theoretical grounding and is potentially inefficient to evaluate or (2) entirely change the MDM framework or architecture, limiting their applicability.
Contribution. In this work, we propose a simple, principled solution addressing these limitations, PRISM: Plug-in Remasking for Inference-time Self-correction of Masked diffusions, a plug-and-play fine-tuning framework that easily adapts to any pretrained MDM (addressing (1)) and directly learns the clean token’s quality given a partially masked sequence (addressing (2)).
Theoretically, PRISM introduces a novel self-correction loss that provably learns the per-token quality scores. It attaches a lightweight adapter to any pretrained MDM and fine-tunes with this loss; at inference, the learned scores detect low-quality tokens to remask and potentially be revised in subsequent inference steps (Figure 1, right). Practically, PRISM improves MDM inference across a range of domains and scales: Sudoku, unconditional text generation with a 170M MDM, and code generation with LLaDA (nie2025large), an open-source 8B MDM. Notably, on LLaDA, with <500<500 GPU-hours of fine-tuning, it acquires the ability to self-correct, outperforming baselines on MBPP austin2021program.
Why self-correction? Although self-correction is a common challenge in generative models where a token cannot be revised, it plays a more critical role in MDMs. This is because of a unique parallel sampling nature of MDMs, which significantly reduces inference cost. However, this parallel generation with only a few inference steps makes them more susceptible to dependency errors across positions (liu2024discrete; xu2024energy; kang2025parallelbench), as MDMs model per-position unmasking posteriors rather than the full joint distribution over positions. This limitation naturally motivates the need for self-correction mechanisms that can detect and rectify erroneous predictions. Indeed, our experiments show that PRISM becomes increasingly effective as more tokens are sampled in parallel, i.e., when MDM inference is performed with fewer sampling steps.
Organization. We begin by reviewing MDMs in Section 2.1 and previous work on equipping MDMs with self-correction ability in Section 2.2. We introduce PRISM in Section 3 with its theoretical foundation and validate its empirical ability in Section 4.
Concurrent work. [^1] This paragraph has been included from version two of this manuscript. Although entailed with different details, (meshchaninov2025guided; li2025riv) address the self-correction problem for MDM by training an additional module that takes $\hat{{\mathbf{x}}}_{0}$ as input, thereby belonging in the class of proxy-$\hat{{\mathbf{x}}}$ approaches in Section 2.2. The most relevant work, huang2025don, shares the same motivation and experimental setup but differs in several aspects, which we elaborate on in Appendix A.
Figure 1: PRISM overview. MDMs learn an unmasking posterior to unmask tokens, after which they remain fixed (Left). PRISM fine-tuning introduces per-token quality, which is used to detect incorrect tokens and remask them. At inference, the fine-tuned MDM jointly computes unmasking posterior and per-token quality, respectively performing unmasking and remasking (Right).
2 Preliminaries
2.1 Masked Diffusion Models
Notation. Assume we aim to learn the distribution over sequences of length $L$ with vocabulary $\mathcal{V}$, e.g., natural language, namely the true distribution ${\mathbf{x}}\sim p_{\mathrm{data}}$. ${\mathbf{x}}^{i}$ denotes the $i$-th element of a given sequence ${\mathbf{x}}=({\mathbf{x}}^{1},\dots,{\mathbf{x}}^{L})$ and $\Delta(\mathcal{V})$ indicates the simplex of probability distributions over $\mathcal{V}$.
MDM training. Although MDMs can be interpreted from several viewpoints, we adopt an any-order language model viewpoint (ou2024your; zheng2024masked), which is essential to understanding our PRISM framework. Roughly speaking, MDM learns the posterior marginals over clean tokens in each masked position, conditioned on a partially masked sequence. We now formalize how this objective is defined and trained. To learn the posterior, MDM performs a masking process using an auxiliary mask token $\mathbf{m}$. For a given ${\mathbf{x}}\sim p_{\mathrm{data}}$, this masking process samples a partially masked sequence ${\mathbf{z}}\in(\mathcal{V}\cup\{\mathbf{m}\})^{L}$ defined in two equivalent ways (see Appendix B.1 for equivalence):
  • (1)
    Draw $n\sim\mathrm{Unif}\{0,\dots,L\}$ and replace the tokens at uniformly selected $n$ indices in $\mathbf{x}$} with $\mathbf{m}$.
  • (2)
    Draw $t\sim\mathrm{Unif}[0,1]$ and for each $i$, replace ${\mathbf{x}}^{i}\in\mathcal{V}$ with $\mathbf{m}$ with probability $t$.
To clarify, in (1), for a given $n$, we uniformly sample an index set from all size-$n$ subsets of $\{1,\dots,L\}$, and mask corresponding positions of $\mathbf{x}$}. This masking process defines a random variable $\mathbf{z}$} conditioned on $\mathbf{x}$}, and marginalizing it over ${\mathbf{x}}\sim p_{\mathrm{data}}$ induces a joint distribution over $({\mathbf{x}},{\mathbf{z}})$.
We refer to the resulting coordinate-wise posterior conditioned on $\mathbf{z}$} as the unmasking posterior; $p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})$, i.e., the distribution over clean tokens at position $i$ given a partially masked sequence $\mathbf{z}$}. This unmasking posterior is the core modeling component in MDM and is parameterized by a neural network $f_{\theta}$, which takes $\mathbf{z}$} as an input and outputs a tensor of shape $|\mathcal{V}|\times L$. Its $i$-th column, $f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})\in\Delta(\mathcal{V})$, models the unmasking posterior $f_{\theta}^{i}(v\,|\,{\mathbf{z}})\approx p({\mathbf{x}}^{i}=v\,|\,{\mathbf{z}})$. To train $f_{\theta}$, MDM minimizes the following loss:
$\mathcal{L}(\theta)\colon=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}}}\left[\frac{1}{n}\sum_{i\colon{\mathbf{z}}^{i}=\mathbf{m}}-\log f_{\theta}^{i}({\mathbf{x}}^{i}\,|\,{\mathbf{z}})\right],$ (1)
which is the cross-entropy loss summed over all masked indices. Here, the expectation is taken over the joint distribution of $({\mathbf{x}},{\mathbf{z}})$ defined above. As desired, the unique minimizer $f_{\theta^{\star}}$ of the loss $\mathcal{L}_{\theta}$ satisfies $f_{\theta^{\star}}^{i}(v\,|\,{\mathbf{z}})=p({\mathbf{x}}^{i}=v\,|\,{\mathbf{z}})$. This connects the MDM training loss to any-order language models such as BERT (devlin2019bert), which model the posterior marginals at all masked positions. We note that the training loss equation 1 allows other equivalent formulations, as developed in previous MDM works (nie2025large; sahoo2024simple; shi2024simplified).
MDM inference. MDM inference starts from a length-$L$ masked sequence ${\mathbf{x}}_{1}=(\mathbf{m},\dots,\mathbf{m})$ and proceeds over a monotonically decreasing sequence of times $t_{0}=1>\dots>t_{N}=0$. At each step $t_{\ell}$, given partially masked sequence ${\mathbf{x}}_{t_{\ell}}\in(\mathcal{V}\cup\{\mathbf{m}\})^{L}$, we proceed by two steps to obtain ${\mathbf{x}}_{t_{\ell+1}}$:
  • (a)
    Choose a subset of masked tokens $\mathcal{S}$ to unmask, $\mathcal{S}\subseteq\{i\,|\,{\mathbf{x}}_{t_{\ell}}^{i}=\mathbf{m}\}$.
  • (b)
    For each $i\in\mathcal{S}$, unmask ${\mathbf{x}}_{t_{\ell}}^{i}$ to a clean token sampled from $f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})\in\Delta(\mathcal{V})$.
Notably, step (a)—the choice of $\mathcal{S}$—is highly flexible and central to MDM’s gains on downstream tasks; $\mathcal{S}$ may be chosen arbitrarily. Two predominant strategies are independently including each masked index ${\mathbf{x}}_{t_{\ell}}^{i}=\mathbf{m}$ with a fixed probability (sahoo2024simple; shi2024simplified) or adaptively selecting indices, e.g., the most confident tokens or the leftmost tokens (chang2022maskgit; zheng2023reparameterized; kim2025train; nie2025large).
2.2 Self-correction for MDM via per-token quality
An ideal generative model would detect low-quality tokens at inference time and correct them when appropriate. Prior work has explored self-correction in LLMs, including kumar2024training; gou2023critic, and, recently, in MDMs. Despite varying instantiations, these methods for MDM can be organized into a common two-step procedure. At each inference step, given ${\mathbf{x}}_{t_{\ell}}$:
  • (a)
    Compute a per-token “quality score” for each ${\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}$ and, according to it, select a subset of clean tokens $\mathcal{T}$ to remask.
  • (b)
    For each $i\in\mathcal{T}$, remask ${\mathbf{x}}_{t_{\ell}}^{i}$ and either replace it with a new clean or leave it masked.
At step (a), the set $\mathcal{T}$ is typically chosen with the lowest per-token quality, optionally with stochasticity. At step (b), the remasked ${\mathbf{x}}_{t_{\ell}}^{i}$ can be immediately resampled (using the pretrained model or an auxiliary module) or kept to engage further for future unmasking/remasking steps. Combined with the standard MDM unmasking step (Section 2.1), inference alternates between unmasking and (re)masking; these may be interleaved or staged, depending on the method.
Challenge. The key challenges are twofold: formulating a principled definition of per-token quality and constructing a model that estimates it. These can be formalized as follows: A per-token quality is a function $g_{\star}$ that takes any partially masked sequence $\mathbf{y}$} and returns position-wise scalars for non-masked positions ${\mathbf{y}}^{i}\neq\mathbf{m}$:
$g_{\star}\colon(\mathcal{V}\cup\{\mathbf{m}\})^{L}\to[0,1]^{L},\quad{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g^{i}_{\star}({\mathbf{y}})\approx(\text{Per-token quality})}=(\text{``how likely'' ${\mathbf{y}}^{i}$ is given ${\mathbf{y}}$}).$
The phrase “how likely” is a modeling choice; whatever the choice, it should align with the intended notion of per-token quality (our PRISM instantiation is given in Section 3). The goal is a computationally efficient algorithm for modeling $g_{\star}$ that, in the limit, provably learns the chosen definition of per-token quality. As we argue next, prior work typically lacks either a precise, testable target of per-token quality or entirely changes the MDM design itself, limiting its applicability. We therefore organize the prior work along whether it specifies a precise target for per-token quality on the state during inference ${\mathbf{x}}_{t_{j}}$ and how disruptive it is to the MDM design.
  • Training-free approaches: DFM (gat2024discrete), ReMDM (wang2025remasking), P2-Self (peng2025pathplanningmaskeddiffusion) are drop-in and can be applied to any pretrained MDM, however they lack a precise target: scores are random (DFM, ReMDM), evaluated on a different input than ${\mathbf{x}}_{t_{\ell}}$ (ReMDM-conf), or inferred from logits that were not trained to present the token quality (P2-Self).
  • Methods that require rehauling MDM design: GIDD (von2025generalized), Informed Corrector (zhao2024informed), indeed articulate a valid target for self-correction but achieve it by changing the pretraining design or architecture, requiring full retraining.
  • Proxy-$\hat{{\mathbf{x}}}$ approaches: Token Critic (gou2023critic) and P2-Train (peng2025pathplanningmaskeddiffusion) train an external scorer to approximate the marginal likelihoods of a sampled clean sequence $\hat{{\mathbf{x}}}$. For a given ${\mathbf{x}}_{t_{\ell}}$, it first samples a reconstructed clean sequence $\hat{{\mathbf{x}}}$ via $f_{\theta}$ in a single step, then calculates the quality score via the trained module. Since the pretrained MDM has learned the per-token unmasking posterior, not the joint posterior across positions, $\hat{{\mathbf{x}}}$ can deviate a lot from a valid sample from the posterior. This method also requires an extra forward pass for scoring.
We provide a comprehensive explanation of the prior work and their notions of per-token quality in Appendix A. In short, prior work either lacks a precise target of per-token quality or alters the design of the MDM itself. In contrast, PRISM assigns an explicit per-token target to a given sequence ${\mathbf{x}}_{t_{\ell}}$ and estimates it in a single forward pass by fine-tuning the pretrained MDM, while still preserving the MDM’s ability to compute per-token posteriors.
3 PRISM
In this section, we introduce Plug-in Remasking for Inference-time Self-correction of Masked diffusion models (PRISM). We substantiate its theoretical foundation in Section 3.1, present our empirical modeling design in Section 3.2, and explain how PRISM is used during inference in Section 3.3. Recall the definition of per-token quality from Section 2.2:
$g_{\star}\colon(\mathcal{V}\cup\{\mathbf{m}\})^{L}\to[0,1]^{L},\quad{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g^{i}_{\star}({\mathbf{y}})\approx(\text{Per-token quality})}=(\text{``how likely'' ${\mathbf{y}}^{i}$ is given ${\mathbf{y}}$}).$ (2)
A natural candidate for the per-token quality is the ground-truth likelihood of the token ${\mathbf{y}}^{i}$ when ${\mathbf{y}}^{i}$ is masked out of $\mathbf{y}$}. We denote this masked sequence by ${\mathbf{y}}\oplus\mathbf{m}_{i}$, i.e., $({\mathbf{y}}\oplus\mathbf{m}_{i})^{i}=\mathbf{m}$ and $({\mathbf{y}}\oplus\mathbf{m}_{i})^{j}={\mathbf{y}}^{j}$ for $j\neq i$. Our desired notion of per-token quality is the likelihood of a given token conditioned on the rest of the sequence. When ${\mathbf{y}}^{i}$ is consistent with the context $\mathbf{y}$}, this likelihood should be high; otherwise, it should be low, marking the position as a candidate for remasking. Formally,
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\star}^{i}({\mathbf{y}})}:=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i}).$
To clarify, $p({\mathbf{x}}^{i}=\cdot\,|\,\cdot)$ is the same as the unmasking posterior used for MDM in Section 2.1.
Discussion on our notion of per-token quality. Since our per-token quality is derived from the posterior, it can be sensitive to uncertainty, e.g., in ${\mathbf{y}}=([\text{I}],\mathbf{m},[\text{to}],[\text{school}])$, although the fourth token $\text{school}]$ is valid, $g_{\star}^{4}({\mathbf{y}})=p({\mathbf{x}}^{4}=[\text{school}]\,|\,([\text{I}],\mathbf{m},[\text{to}],\mathbf{m}))$ may be low since there are multiple candidates for that position. In such situations, it cannot “hurt” to remask/resample such tokens, as there is high entropy in the true posterior marginal anyway.
More importantly, our notion of per-token quality is effective at detecting incorrect tokens in positions with low uncertainty, e.g., in ${\mathbf{y}}=([\text{I}],[\text{eat}],[\text{a}],[\text{apple}],[\text{strudel}])$, the third token $\text{a}]$ is ungrammatical given $\text{apple}]$, and $g_{\star}^{3}({\mathbf{y}})$ is sharply concentrated on other words like $\text{an}]$$\text{the}]$; our quality score would correctly identify $\text{a}]$ for remasking and revision.
Challenge: Unmasking vs. Remasking training. Our goal is to train (or obtain) a model $g_{\phi}$ that approximates $g_{\star}$. This is challenging because $g_{\phi}$ must take the sequence where the position $i$ is already unmasked, yet it must approximate the posterior defined with the masked context ${\mathbf{y}}\oplus\mathbf{m}_{i}$. In contrast, an MDM $f_{\theta}$ predicts from the masked position itself;
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})\in[0,1],\quad{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})}\approx p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})\in\Delta(\mathcal{V}),$
Although $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})$} and the $\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$}’s ${\mathbf{y}}^{i}$-th coordinate target the same quantity, they condition on different inputs. Consequently, $g_{\phi}$ cannot be simply trained or efficiently queried as in MDM.
A naive workaround is a distillation-based approach, supervising $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})$} with the teacher signal $\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{y}}\oplus\mathbf{m}^{i})$}. This has two drawbacks: First, even in an infinite data and compute regime, we cannot achieve perfect $g_{\phi^{\star}}$ if the teacher model $f_{\theta}$ is imperfect. Second, it is data-inefficient: each training pair $({\mathbf{y}},i)$ requires a fresh evaluation of $f_{\theta}$ on ${\mathbf{y}}\oplus\mathbf{m}_{i}$. Consequently, one forward pass of $f$ supervises only a single index, i.e., covering $m$ clean token positions in a sequence requires $m$ passes.
To address this input mismatch, zhao2024informed leverage a hollow-transformer $g_{\phi^{\prime}}$, with the property $g_{\phi^{\prime}}^{i}({\mathbf{y}}\oplus\mathbf{m}_{i})\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$, enabling the use of the same loss as MDM. However, this requires adopting a new architecture (with a tweak on the attention mechanism) that processes $\mathbf{y}$} while equivalently behaving as if position $i$ were masked, which entails retraining and limiting plug-and-play applicability to pretrained MDMs.
As we will show, PRISM does not possess these limitations: it is plug-and-play with any pretrained MDM $f_{\theta}$ (no architectural changes), and—even when $f_{\theta}$ is imperfect, it recovers the desired per-token quality $g_{\phi}$ in the infinite-data limit.
3.1 Theoretical foundation of PRISM
In this section, we establish the theoretical foundation of PRISM. We begin by recalling the premise behind the MDM training loss before describing how we adapt this premise to give rise to PRISM.
Marginalization perspective on MDM training. We rewrite the MDM training loss in equation 1.
$\mathcal{L}(\theta)\coloneqq\mathbb{E}_{{\mathbf{x}},{\mathbf{z}}}\left[\frac{1}{n}\sum_{i\colon{\mathbf{z}}^{i}=\mathbf{m}}-\log f_{\theta}^{i}({\mathbf{x}}^{i}\,|\,{\mathbf{z}})\right].$
We now aim to understand why the $\mathcal{L}(\theta)$ has the unmasking posterior as its unique minimizer. First, the expectation over the joint distribution of $({\mathbf{x}},{\mathbf{z}})$ can be written by first sampling $\mathbf{z}$} and then $\mathbf{x}$} from the induced posterior. For fixed $\mathbf{z}$}, the loss only depends on the posterior of ${\mathbf{x}}^{i}$ given $\mathbf{z}$}, which is exactly the unmasking posterior $p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})$. Thus, we can rewrite:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{z}}\left[\frac{1}{n}\sum_{i\colon{\mathbf{z}}^{i}=\mathbf{m}}{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}\mathbb{E}_{v\sim p({\mathbf{x}}^{i}=\cdot\mid{\mathbf{z}})}\left[-\log f_{\theta}^{i}(v\,|\,{\mathbf{z}})\right]}\right],$
and then we observe that the blue-colored term is the cross-entropy between two distributions $f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})$ and $p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})$, which is minimized at $f_{\theta^{\star}}^{i}(\cdot\mid{\mathbf{z}})=p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})$ for every $({\mathbf{z}},i)$. Put differently, the masking process of MDM defines the joint distribution over $({\mathbf{x}},{\mathbf{z}})$ and fixing $\mathbf{z}$} and marginalizing over $\mathbf{x}$} yields a cross-entropy between the network output and the desired unmasking posterior. We use this marginalization technique to build PRISM.
PRISM. Recall our goal for training $g_{\phi}$ to predict per-token quality, formally defined as $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi^{\star}}^{i}({\mathbf{y}})=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$}. Given a pretrained MDM $f_{\theta}$, PRISM proceeds as follows.
  • (a)
    Construct a pair $({\mathbf{x}},{\mathbf{z}})$ as in MDM training: sample ${\mathbf{x}}\sim p_{\mathrm{data}}$ and then draw $\mathbf{z}$} by masking.
  • (b)
    Sample $\mathbf{y}$}: choose a masked index $i$ from $\mathbf{z}$} and replace it with a clean token ${\mathbf{y}}^{i}\sim f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})$.
Step (a) defines the same joint distribution over $({\mathbf{x}},{\mathbf{z}})$ as MDM. Step (b) yields a random variable $({\mathbf{y}},i)$ conditioned on $\mathbf{z}$}. Although we do not designate a sampling rule for the index $i$, PRISM’s guarantee is invariant to its rule, e.g., uniformly random, confidence-based. As a result, PRISM induces a joint distribution over $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$. Taking expectation over this joint distribution, we define the PRISM loss:
$\mathcal{L}(\phi)\coloneqq\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\left[\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}\,)\right],$ (3)
where $\mathrm{BCE}$ denotes a Binary Cross-Entropy loss $\mathrm{BCE}(b,p)\!=\!-b\log p\!-\!(1\!-\!b)\log(1\!-\!p)$ for a binary label $b\!\in\!\{0,1\}$. This loss is the crux of PRISM, whose unique minimizer is the true per-token quality.
Proposition 1 (PRISM).
The per-token quality uniquely minimizes the PRISM loss $\mathcal{L}(\phi)$ equation 3: $g_{\phi^{\star}}^{i}({\mathbf{y}})=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i}).$
Before proving Proposition 1, we highlight two key advantages that PRISM offers.
Advantages. First, the minimizer $g_{\phi^{\star}}^{i}({\mathbf{y}})$ does not depend on the specific choice of $f_{\theta}$; $f_{\theta}$ affects the training objective only through the marginalized distribution of $({\mathbf{y}},i)$, not the minimizer. This contrasts with the aforementioned distillation-based approach, in which an imperfect $f_{\theta}$ will not give rise to a perfect $g_{\phi}$. Second, PRISM naturally adapts to a pretrained MDM; a single model can share the same architecture and carry two heads, one for the unmasking posterior and one for per-token quality. We detail the modeling choices of PRISM in Section 3.2.
PRISM: Guarantee. Proposition 1 can be demonstrated within a few lines. The key mechanism is that the joint distribution $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$ exhibits a simple posterior given $({\mathbf{y}},i)$; $\mathbf{z}$} is deterministically ${\mathbf{z}}={\mathbf{y}}\oplus\mathbf{m}_{i}$ and each ${\mathbf{x}}^{i}$ is drawn from the same unmasking posterior of MDM, i.e., $v\sim p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})$. Therefore, fixing $({\mathbf{y}},i)$ and then marginalizing the binary label $\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]$ over $({\mathbf{x}},{\mathbf{z}})$ yields the simple byproduct,
${\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}q}\coloneqq\mathbb{E}_{v\sim p({\mathbf{x}}^{i}=\cdot\mid{\mathbf{y}}\oplus\mathbf{m}_{i})}\left[\mathbf{1}[v={\mathbf{y}}^{i}]\right]=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i}),$
the per-token quality. We then expand the binary cross-entropy loss:
$\mathcal{L}(\phi)=\mathbb{E}_{({\mathbf{y}},i)}\left[-{\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}q}\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})})-({\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}1-q})\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}1-g_{\phi}^{i}({\mathbf{y}})})\right],$
which is exactly minimized at $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi^{\star}}^{i}({\mathbf{y}})=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$}.
3.2 Empirical design choice of PRISM
In this section, we present our empirical modeling choices that go into the design of PRISM. We first explain the plug-and-play aspect of PRISM, in which we adapt a pretrained MDM and fine-tune it via the PRISM loss.
Figure 2: PRISM training pipeline. Fine-tuning samples are built in two steps; (a) masking process to obtain $({\mathbf{x}},{\mathbf{z}})$ and (b) unmask a few indices in $\mathbf{z}$} to get $\mathbf{y}$} with a pretrained MDM $f_{\theta}$ (Left). We add a lightweight adapter to a pretrained MDM so that the resulting model computes the unmasking posterior ($f_{\theta}$) and per-token quality ($g_{\theta}$) (Right).
Plug-and-Play approach. Assume a pretrained MDM $f_{\theta}$ is given, which by design has learned the unmasking posterior. PRISM augments this model with a lightweight auxiliary adapter that reuses the same backbone to produce a size-$L$ tensor. In words, the unmasking posterior and per-token quality share a single backbone and are computed by separate heads (Figure 2).
For a given sequence, we respectively denote the two head outputs as $f_{\theta}$ and $g_{\theta}$. Let $\mathbf{h}_{\theta}$ denote the final hidden state of the MDM backbone. The unmasking posterior $f_{\theta}$ is obtained by passing $\mathbf{h}_{\theta}$ through the unmasking head (which originally exists in pretrained MDM) and then applying the softmax. The per-token quality $g_{\theta}$ is obtained by passing $\mathbf{h}_{\theta}$ through the attached head and applying the sigmoid. Sigmoid is used to guarantee $g_{\theta}^{i}\in[0,1]$, making it suitable for the binary cross-entropy loss. Thus, with a single $\theta$ and a lightweight adapter, we jointly model the unmasking posterior and then the per-token quality: for a partially masked sequence $\mathbf{w}$},
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{w}})}\approx p({\mathbf{x}}^{i}={\mathbf{w}}^{i}\,|\,{\mathbf{w}}\oplus\mathbf{m}_{i})\in[0,1],\quad{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{w}})}\approx p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{w}})\in\Delta(\mathcal{V}).$
At a high level, the PRISM recipe is simple: train $g_{\theta}$ using the PRISM loss defined in equation 3. Consequently, the parameters updated through $g_{\theta}$ are those of the attached head and the backbone, excluding the unmasking head. As an alternative to fine-tuning the backbone, one can add a LoRA adapter (hu2022lora) into the backbone and train only the LoRA adapter parameters.
Practical interventions. We introduce two practical interventions that improve PRISM’s efficiency. First, to preserve $f_{\theta}$’s ability to estimate the unmasking posterior while $g_{\theta}$ is learning toward the per-token quality, we add the MDM training loss equation 3 as a regularization term. For a regularization constant $\lambda>0$, the total loss is therefore:
$\mathcal{L}(\theta)=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\Bigg[{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{y}})}\,)}+\lambda\times\frac{1}{n}\sum_{j\colon{\mathbf{z}}^{j}=\mathbf{m}}{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}-\log f_{\theta}^{j}({\mathbf{x}}^{j}\,|\,{\mathbf{z}})}\Bigg].$ (4)
Second, in the sampling process for constructing PRISM training pairs $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$, we modify step (b). To obtain $\mathbf{y}$}, rather than unmasking a single $\mathbf{y}$} from $\mathbf{z}$}, we obtain multiple $({\mathbf{y}},i)$ pairs from one $\mathbf{z}$}. This produces multiple training pairs $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$ with varying $({\mathbf{y}},i)$ values from a single forward pass of $f_{\theta}(\cdot\,|\,{\mathbf{z}})$, improving data efficiency. Since this procedure of sample construction minimizes $g_{\theta}$ (not $f_{\theta}$), we apply stop-gradient to this $f_{\theta}$. The pseudocode is given as Algorithm 1.
Importantly, the theoretical guarantee in Proposition 1 still maintains: the true minimizer $g_{\theta}$ of this modified PRISM loss continues to encode meaningful per-token quality. We defer the precise statement and proof to Appendix B.2. Moreover, we conduct a careful ablation of different design choices on sampling $\mathbf{y}$} in Appendix D.3 and Appendix D.4.
Practical advantages. PRISM formulation not only guarantees to minimize the ground truth per-token quality but also has two practical advantages. First, since $f_{\theta}$ is already pretrained, its hidden states encode useful representations, likely accelerating $g_{\theta}$’s training on the per-token quality. Second, the targeted per-token quality is a scalar per position, in contrast to the per-position distribution by the unmasking posterior, making it substantially easier to learn. Perhaps surprisingly, we observe that using much less compute compared to MDM pretraining already yields a meaningful learning of the per-token quality (which we detail these results in Section 4), reinforcing the perspective that PRISM is a cheap fine-tuning intervention.
Algorithm 1 PRISM fine-tuning from a pretrained MDM
1:Require: Pretrained MDM backbone $f_{\theta}$, per-token quality head $g_{\theta}$, optimizer $\mathcal{O}$, true distribution $p_{\mathrm{data}}$, regularization constant $\lambda>0$, number of unmasking tokens $k\in\mathbb{N}$.
2:Sample ${\mathbf{x}}\sim p_{data}$, $n$, and $\mathbf{z}$} from the masking process.
3:Initialize the loss $\mathcal{L}(\theta)\leftarrow 0$
[…]
5:  # Sample $\mathbf{y}$} from $\mathbf{z}$} with $f_{\mathrm{sg}(\theta)}$
6:  Select a subset $\mathcal{S}\subseteq\{i\mid{\mathbf{z}}^{i}=\mathbf{m}\}$ with $|\mathcal{S}|=k$.
7:  Obtain $\mathbf{y}$} by replacing ${\mathbf{z}}^{i}=\mathbf{m}$ to $v^{i}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f_{\mathrm{sg}(\theta)}^{i}(\cdot\,|\,{\mathbf{z}})}$ for each $i\in\mathcal{S}$.
8:  # Calculate Loss
9:  $\mathcal{L}(\theta)\leftarrow\mathcal{L}(\theta)+\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{y}})}}\,)}$ $\triangleright$ PRISM Loss
10:  $\mathcal{L}(\theta)\leftarrow\mathcal{L}(\theta)+\lambda\times\frac{1}{n}\sum_{j:{\mathbf{z}}^{j}=\mathbf{m}}{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}-\log f_{\theta}^{j}({\mathbf{x}}^{j}\,|\,{\mathbf{z}})}$. $\triangleright$ Regularization (MDM loss)
11:  $\theta\leftarrow\mathcal{O}(\theta,\mathcal{L}(\theta))$
12:end while
3.3 Employing PRISM at inference
In this section, we explain how a model resulting from PRISM is used to perform remasking at inference. The high-level procedure follows Section 2.1; at each inference step, we perform remasking and unmasking simultaneously, using the learned unmasking posterior and per-token quality, respectively. Assume we are given a fine-tuned MDM $\theta$ (via PRISM loss) and a monotonically decreasing grid of times $t_{0}=1>\dots>t_{N}=0$ for inference. We initialize ${\mathbf{x}}_{t_{0}}=(\mathbf{m},\dots,\mathbf{m})$. At each step $t_{\ell}$, we obtain ${\mathbf{x}}_{t_{\ell+1}}$ by these following steps.
  • (a)
    Choose a subset of masked tokens $\mathcal{S}$ to unmask.
  • (b)
    Choose a subset of clean tokens $\mathcal{T}$ to remask with the lowest quality scores $g_{\theta}({\mathbf{x}}_{t_{\ell}})$.
  • (c)
    For each $i\in\mathcal{T}$, remask ${\mathbf{x}}_{t_{\ell}}^{i}$ and for each $j\in\mathcal{S}$, unmask ${\mathbf{x}}_{t_{\ell}}^{j}$ to a clean token sampled from $f_{\theta}^{j}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})\in\Delta(\mathcal{V})$.
We emphasize that $f_{\theta}^{\cdot}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})$ and $g_{\theta}({\mathbf{x}}_{t_{\ell}})$ are obtained from a single forward pass of the backbone. Consequently, our method does not add any computational overhead relative to methods without remasking (vanilla MDM inference) or training-free approaches, e.g, ReMDM.
[…]
4 Experiments
In this section, we demonstrate that PRISM is a practically efficient and scalable approach across different domains and scales. In Section 4.1, we apply PRISM to a 3030M-scale MDM for Sudoku. In Section 4.2, we apply PRISM to a 170M-scale MDM and evaluate its ability for natural language modeling. Finally, in Section 4.3, we apply PRISM to LLaDA-8B-Instruct (nie2025large) and evaluate it on a coding task. We first outline the pipeline shared across all experiments. We provide the experimental details in Appendix D.
Experiment pipeline. We start from an MDM pretrained for each domain and attach a lightweight adapter. We then fine-tune the model on the PRISM loss defined in equation 4. For baselines, we adopt ReMDM (ReMDM-cap or ReMDM-loop) and ReMDM-conf from wang2025remasking; for fair comparison, we re-evaluate these baselines using the same backbone after PRISM fine-tuning. Finally, since our focus is on the remasking strategy, we restrict comparisons to remasking baselines. We summarize our experimental results as follows:
  • Efficiency: The PRISM fine-tuned MDM has the ability to self-correct using much fewer samples compared to those used for pretraining.
  • Performance: The PRISM-tuned MDMs outperform ReMDM and ReMDM-conf, indicating that PRISM learns our desired notion of per-token quality quite well.
4.1 Motivating example: Sudoku puzzle
In this section, we examine PRISM on Sudoku puzzles and show that the learned per-token quality operates as desired. Before moving on to the results, we use Sudoku as a metaphorical testbed to explain why ReMDM and ReMDM-conf can fail to capture the intended per-token quality.
Sudoku as a metaphorical testbed. ReMDM assigns token quality randomly, which can unnecessarily flag correct cells. ReMDM-conf sets the token quality to the likelihood at the time a cell was unmasked; this is computed on a previous Sudoku board state and can quickly become misaligned from the current board. In contrast, PRISM predicts per-token quality under the current board, thereby reliably identifying incorrect cells.
Setup. We construct a corpus of Sudoku puzzles (4848k for training and 22k for evaluation), following the setup of (ben2025accelerated; alp2024sudoku). Each board is tokenized as a length-$89$ sequence comprising $81$ cell values and $8$ [EOL] tokens that separate columns. Following the MDM pretraining recipe of sahoo2024simple, we use a 28.628.6M Diffusion Transformer (DiT; peebles2023scalable) architecture as the backbone and pretrain it for 100100k iterations ($\approx 530$ epochs). We then fine-tune it with the PRISM loss, additionally with a lightweight adapter (0.130.13M parameters).
Figure 3: Left: PRISM achieves higher success accuracy on Sudoku puzzles than baselines. (red line) Right: PRISM detects the incorrect cells by assigning low per-token quality (red-colored cells).
Results. We evaluate the success rate for the same model across epochs for PRISM and baselines. As shown in Figure 3, PRISM achieves progressively higher success rates during fine-tuning (red line) and surpasses the baselines. This confirms that the learned per-token quality enables effective self-correction at inference. Notably, PRISM starts to outperform vanilla MDM inference (dashed line) and baselines within only a few epochs, far fewer than the pretraining epochs ($\approx 530$), corroborating our claim of sample efficiency.
Visualization of learned per-token quality. Figure 3 further visualizes PRISM’s ability to detect incorrect tokens. On the shown board, three digits are misfilled: $(2,6)=8$, $(5,5)=4$, and $(7,3)=7$. The computed per-token quality assigns low scores to these cells (red), while other cells attain high scores, nearly $0.99$ (blue), indicating that the fine-tuned model has almost perfectly learned the per-token quality.
4.2 Unconditional Text Generation
Setup. We first take a DiT-based MDM with 170170M parameters, using a checkpoint from (sahoo2024simple). This MDM is pretrained on the OpenWebText corpus (Gokaslan2019OpenWeb) with the GPT-2 tokenizer (radford2019language), a maximum sequence length $L=1024$, and a total of 262262B tokens. We then attach a lightweight 77M parameter adapter, one attention layer applied to the final hidden state, and fine-tune it with the PRISM loss on a total 164164M tokens from the same OpenWebText corpus. This fine-tuning is data-efficient, using $\mathbf{1600}\times$ fewer tokens than pretraining.
Results. We then assess the performance of PRISM and baselines on unconditional generation. Our primary metrics are Generative Perplexity and MAUVE (pillutla2021mauve), which respectively assess per-sequence fluency and distributional similarity between generated samples and the reference distribution. Moreover, given the prior observation (zheng2024masked) that a sequence with too many redundant tokens can spuriously yield better (i.e., lower) generative perplexity, we also report the entropy as a sanity metric. Following standard evaluation protocols (pillutla2021mauve; wang2025remasking), we generate $5000$ samples to compute both metrics, using GPT-2-Large as the reference model. As shown in Figure 4, PRISM outperforms baselines on both metrics (red lines) while achieving a comparable entropy, especially in the regime with fewer sampling steps ($N<1024$).
Sampling strategy. As sentence diversity is crucial for unconditional generation, we observe that the vanilla PRISM inference from Section 3.3 can degrade diversity. We attribute this to the greedy nature of the token-remasking policy that ranks positions by per-token quality. To mitigate this, we adopt PRISM-loop, analogous to ReMDM-loop (wang2025remasking): follow PRISM’s remasking scheme and add loop steps at $t_{N}=0$. After obtaining a clean sequence, in a loop, we randomly remask a subset of tokens and unmask them. A full breakdown of the sampling strategy is given in Appendix C.1,  D.2.
Figure 4: Assessing PRISM on unconditional text generation. PRISM (red) exhibits superior performance compared to baselines (ReMDM: green, ReMDM-conf: blue, Vanilla MDM: gray), particularly in the regime with fewer sampling steps ($N<1024$).
4.3 Appyling PRISM to 88B MDM
Setup. We take LLaDA-88B-Instruct (nie2025large), an open-sourced 88B MDM. While freezing the backbone, we add an additional head and a LoRA adapter (hu2022lora), for a total of $\approx 250$M trainable parameters. We fine-tune it using the PRISM loss on the opc-sft-stage2 (Huang2024OpenCoderTO) (0.10.1M pairs) for $100$ epochs. Fine-tuning completes in $\approx 30$ hours on $12$ H100 GPUs.
Results. We evaluate the resulting model in a zero-shot setting on MBPP (austin2021program), a challenging Python coding benchmark with $974$ problems. As shown in Table 1, PRISM outperforms ReMDM and ReMDM-conf, especially in the regime of $256$ sampling steps, by approximately $2.7\%$. We detail training and inference configurations in Appendix D.5D.6.
Sampling steps 256 512 1024
PRISM 23.0% 30.5% 32.9%
ReMDM 20.2% 29.7% 33.2%
ReMDM-conf 20.3% 30.0% 32.7%
MDM 18.6% 28.2% 32.7%
Table 1: PRISM on LLaDA-8B-Instruct.
5 Conclusion
Discussion on experimental results. MDMs model per-position unmasking posteriors rather than the joint distribution over positions. Consequently, with few-step inference, they are prone to dependency errors across positions (liu2024discrete; xu2024energy), which creates a natural opportunity for our self-correction mechanism to detect erroneous predictions. We confirm this empirically: methods that incorporate remasking outperform vanilla MDM inference as the number of sampling steps decreases (Figure 4, Table 1). In particular, PRISM achieves larger gains than baselines by learning per-token quality. However, as the number of sampling steps increases, partially unmasked sequences contain fewer errors, leaving less headroom for the self-correction. For the scaling-up experiments in Section 4.3, although we fine-tune LLaDA-8B-Instruct on only 0.1M dataset pairs and evaluate it on MBPP, we expect that a more generalized model could be obtained by constructing a richer and more diverse dataset, given sufficient compute.
Outlook. We introduced PRISM, a plug-and-play fine-tuning framework that equips pretrained MDMs with self-correction ability, addressing prevalent limitations of prior work. PRISM operates under a theoretically grounded fine-tuning loss and shows efficiency across multiple domains and scales.
A unique feature of the PRISM loss is that it provably learns per-token quality without fully generating a clean sequence. This sheds light on PRISM’s potential compared to other approaches, such as reinforcement learning, where full sequence generation is necessary to obtain a learning signal. That said, our notion of per-token quality score has limits; as it is based on the per-position posterior marginals, it cannot fully capture global errors, e.g., reasoning. An important direction for future work is to develop frameworks that enable discrete diffusion models to detect global reasoning errors.
Stepping back, this work is part of a broader program to build generative models that compose discrete sequences the way that humans do–through correction and reordering. The flexibility of MDMs at inference time offers a key stepping stone towards this goal.
Appendix A Related Works
Masked Diffusion Models. Masked Diffusion Models (MDMs) were previously built on continuous-time Markov chains over discrete spaces (hoogeboom2021argmax). Subsequently, austin2021structured introduced D3PM with several families of transition kernels, followed by SEDD (lou2023discrete). These discrete diffusions can also be understood through the lens of discrete flows (campbell2024generative; gat2024discrete), which parallel flow matching and stochastic interpolants over continuous spaces (albergo2022building; albergo2023stochastic; lipman2022flow). Among transition-kernel designs, the absorbing state (masking) kernel has been particularly popular due to its simple formulation and training objective. Subsequent work established a theoretical framework for these models (zheng2023reparameterized; sahoo2024simple; zheng2024masked). Follow-up works have scaled MDMs up to billions of parameters, showcasing their potential compared to autoregressive models (nie2025large; dream2025; gemini2025diffusion; labs2025mercury; gong2025diffucoder).
Inference-time strategies for discrete diffusion models. A key feature of MDMs is their inference-time flexibility, highlighted by zheng2024masked; ou2024your and aligned with our any-order perspective in Section 2.1. Recall the MDM inference procedure from Section 2.1:
  • (a)
    Choose a subset of masked positions $\mathcal{S}\subseteq\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}=\mathbf{m}\}$ to unmask.
  • (b)
    For each $i\in\mathcal{S}$, unmask by sampling a clean token from $f_{\theta}^{i}(\cdot\mid{\mathbf{x}}_{t_{\ell}})\in\Delta(\mathcal{V})$ and setting ${\mathbf{x}}_{t_{\ell}}^{i}$ accordingly.
The choice of $\mathcal{S}$ is entirely flexible, and an appropriate selection can substantially improve downstream task performance. Since our work focuses on self-correction, i.e., re-masking, we only briefly survey prior work on unmasking strategies (kim2025train; peng2025pathplanningmaskeddiffusion; chang2022maskgit; ben2025accelerated; rout2025anchored). A complementary line of work studies inference-time planning strategies for discrete diffusion with uniform transition kernels (liu2024think; varma2024glauber); these methods fall outside the scope of MDMs.
Comparison to huang2025don. For clarity, we provide a comprehensive comparison with huang2025don. The overall training pipeline is largely similar but differs in several key aspects: (1) we provide theoretical guarantees for our training method (PRISM), and (2) we use a pretrained MDM to sample from $\mathbf{z}$} to $\mathbf{y}$}, whereas huang2025don construct $\mathbf{y}$} via a random corruption process. Nevertheless, their large-scale experiments on LLaDA-8B-Instruct cover a broader range of tasks and datasets, while our work focuses on relatively smaller-scale experiments.
A.1 Details on prior work on remasking strategy (self-correction) for MDM
In this section, we detail our discussion in Section 2.2, focusing on their different notions of per-token qualities. We recall the notion of per-token quality stated in Section 2.2.
$g_{\star}\colon(\mathcal{V}\cup\{\mathbf{m}\})^{L}\to[0,1]^{L},\quad{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g^{i}_{\star}({\mathbf{y}})\approx(\text{Per-token quality})}=(\text{``how likely'' ${\mathbf{y}}^{i}$ is given ${\mathbf{y}}$}).$
  • DFM (gat2024discrete) and ReMDM (wang2025remasking) randomly remask clean tokens in $\mathbf{y}$}, which is equivalent to assigning i.i.d. per-token qualities $g_{\star}^{i}({\mathbf{y}})\sim\operatorname{Unif}[0,1]$.
  • P2-Self (peng2025pathplanningmaskeddiffusion) defines per-token quality using the pretrained MDM’s score for the observed token under input $\mathbf{y}$}, e.g., the probability $f_{\theta}^{i}({\mathbf{y}}^{i}\mid{\mathbf{y}})$. Given that ${\mathbf{y}}^{i}\neq\mathbf{m}$, the model was never trained on this index in this state; it only encodes a meaningful quantity (unmasking posterior) when ${\mathbf{y}}^{i}=\mathbf{m}$. This notion of per-token quality is therefore not grounded.
  • ReMDM-conf (wang2025remasking) takes the per-token quality to be the likelihood at the step the token was unmasked: if position $i$ was unmasked at $t_{k}>t_{\ell}$ (i.e., an earlier step), set $g_{\star}^{i}:=f_{\theta}^{i}({\mathbf{x}}_{t_{k}}^{i}\mid{\mathbf{x}}_{t_{k}})=f_{\theta}^{i}({\mathbf{x}}_{t_{\ell}}^{i}\mid{\mathbf{x}}_{t_{k}})$. As noted, this ignores the current state ${\mathbf{x}}_{t_{\ell}}$ and can therefore be misaligned with the model’s present beliefs.
  • Token Critic (lezama2022improved; lezama2022discrete) and P2-train (peng2025pathplanningmaskeddiffusion) train an auxiliary scorer for the likelihood of a given prediction of clean sequence $\hat{{\mathbf{x}}}$. Given ${\mathbf{x}}_{t_{\ell}}$, they first form a full completion $\hat{{\mathbf{x}}}$ by unmasking each mask token at position $i$ from $f_{\theta}^{i}(\cdot\,\mid\,{\mathbf{x}}_{t_{\ell}})$, then evaluate quality by passing this $\hat{{\mathbf{x}}}$ to the external module. A key limitation is that $\hat{{\mathbf{x}}}$ is drawn from per-token unmasking posteriors, not the ground-truth posterior across positions; thus, even with a perfect pretrained MDM, $\hat{{\mathbf{s}}}$ can deviate substantially from the true likelihood ${\mathbf{x}}\sim p_{\mathrm{data}}$, losing a theoretical guarantee.
  • GIDD (von2025generalized) introduces a family of transition kernels that linearly combine uniform and absorbing (masking) kernels. Models pretrained under this kernel encode posteriors at clean-token positions–however, this requires training under the new kernel and is not directly applicable to an off-the-shelf pretrained MDM, altogether losing the benefit of MDM over other types of kernels.
  • Informed corrector (zhao2024informed) employs a hollow-transformer $g_{\phi^{\prime}}$ satisfying $g_{\phi^{\prime}}^{i}({\mathbf{y}})=g_{\phi^{\prime}}^{i}({\mathbf{y}}\oplus\mathbf{m}_{i})\approx p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\mid{\mathbf{y}}\oplus\mathbf{m}_{i})$, enabling the same training loss as MDM. This, however, mandates a modified architecture (attention tweak) that processes $\mathbf{y}$} as if position $i$ were masked, necessitating retraining.
In summary, while these methods improve over vanilla MDM inference without remasking, they either substantially reshape the MDM pipeline or rely on imprecise notions and proxies for per-token quality.
Appendix B Training algorithms and guarantees of PRISM
B.1 Marginal Equivalence of the Masking Process
We formalize that the two masking procedures stated in Section 2.1 induce the same conditional distribution of $\mathbf{z}$} given $\mathbf{x}$}, hence the same joint distributions $({\mathbf{x}},{\mathbf{z}})$. It suffices to show they generate the same distribution over masked index sets.
Consider (b): draw $t\sim\mathrm{Unif}[0,1]$ and for each $i$, replace ${\mathbf{x}}^{i}\in\mathcal{V}$ with $\mathbf{m}$ with probability $t$. The probability of $n$ tokens to be masked is:
$\int_{0}^{1}\binom{L}{n}t^{n}(1-t)^{L-n}dt=\binom{L}{n}\mathrm{B}(n+1,\,L-n+1)=\frac{1}{L+1},$
where $\mathrm{B}$ denotes Beta function. This coincides with procedure (a), where $n$ is uniformly sampled from $\{0,1,\dots,L\}$.
B.2 PRISM provably learns the per-token quality
In this section, we restate Proposition 1 and detail its proof. The proof essentially follows Section 3.1, but we provide the detailed calculations here.
Proposition 1 (restated). Let the PRISM loss be
$\mathcal{L}(\phi)=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\!\left[\mathrm{BCE}\big(\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}],\,g_{\phi}^{i}({\mathbf{y}})\big)\right],\quad\mathrm{BCE}(b,p)=-\,b\log p-(1-b)\log(1-p).$
Then the unique minimizer satisfies, for every $i$ and $\mathbf{y}$},
$g_{\phi^{\star}}^{i}({\mathbf{y}})\;=\;p\!\left({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,\middle|\,{\mathbf{y}}\oplus\mathbf{m}_{i}\right).$
Proof.
First, the expectation over the joint distribution $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$ can be written by first sampling $({\mathbf{y}},i)$ and then $({\mathbf{z}},{\mathbf{x}})$ from the induced posterior. We observe that this induced posterior admits a handy expression:
${\mathbf{z}}={\mathbf{y}}\oplus\mathbf{m}_{i},\quad v\sim p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}}).$
Here $p({\mathbf{x}}^{i}=\cdot\,|\,\cdot)$ is the unmasking posterior given by the MDM’s masking process (Section 3.1), and to avoid notational confusion, we introduce a dummy variable $v$. Next, we expand the expectation term.
$\mathcal{L}(\phi)=$ $\mathbb{E}_{({\mathbf{y}},i)}\left[\mathbb{E}_{v\sim p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}}),z={\mathbf{y}}\oplus\mathbf{m}_{i}}\left[\mathrm{BCE}\big(\mathbf{1}[v={\mathbf{y}}^{i}],\,{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}\big)\right]\right]$
$=$ $\mathbb{E}_{({\mathbf{y}},i)}\left[-{\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}q}\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})})-({\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}1-q})\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}1-g_{\phi}^{i}({\mathbf{y}})})\right],$
where
${\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}q}\coloneqq\mathbb{E}_{v\sim p({\mathbf{x}}^{i}=\cdot\mid{\mathbf{y}}\oplus\mathbf{m}_{i})}\left[\mathbf{1}[v={\mathbf{y}}^{i}]\right]=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i}).$
Finally, we conclude that $g_{\phi^{\star}}^{i}({\mathbf{y}})=q=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$ since $f_{q}(x)=-q\log x-(1-q)\log(1-x)$ is minimized at $x=q\in(0,1)$. ∎
B.3 Extension on PRISM’s provable guarantee
In this section, we provide the pseudocode of the practically used PRISM fine-tuning pipeline and then state its theoretical guarantee.
We note that the choice of index set $\mathcal{S}$ at line 5 of Algorithm 1 is flexible; it can either be random or confidence-based. To clarify, at line 6, although we use $f_{\theta}$ to obtain $\mathbf{y}$}, we use the stop-gradient operation.
Note that the process of obtaining $\mathbf{y}$} and $\mathcal{S}$ can be interpreted as a process of obtaining multiple $({\mathbf{y}},i)$ pairs for each $i\in\mathcal{S}$. Under this, the expectation over $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},\mathcal{S}))$ is equivalent to the expectation over $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$. Therefore, both formulations of the loss function are valid and equivalent.
$\mathcal{L}(\theta)=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},\mathcal{S})}\left[\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{y}})}}\,)}\right]=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\left[{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{y}})}}\,)}\right].$
Now we provide the theoretical guarantee on our PRISM loss.
Proposition 2 (Provable guarantee of PRISM-extension).
For a fixed $1\leq k\leq L$, the PRISM loss
$\mathcal{L}(\phi)=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\left[{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}}\,)}\right]$
has the unique minimizer
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi^{\star}}^{i}({\mathbf{y}})}=\mathbb{E}_{\mathcal{S}\sim\Pi(\cdot\mid{\mathbf{y}},i)}\left[p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\mid{\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}})\right],$
where the expectation is taken over the posterior distribution over $\mathcal{S}$ given $\mathbf{y}$} and $i$, induced by the joint distribution above. Here ${\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}}$ is a sequence obtained from $\mathbf{y}$} by replacing its $i$-th token with $\mathbf{m}$ for every $i\in\mathcal{S}$, aligning the notation in Section 3.1.
Proof.
Before proving this, we clarify some notations. The notation ${\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}}$ extends the notation ${\mathbf{y}}\oplus\mathbf{m}_{i}$ from Section 3.1 into the case where we mask multiple tokens, more precisely, ${\mathbf{y}}\oplus\mathbf{m}_{\{i\}}$.
[…]
By using this property, we now complete the proof; the remaining parts are the same as Proposition 1.
$\mathcal{L}(\phi)=$ $\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\left[{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}}\,)}\right]$
$=$ $\mathbb{E}_{y,i}\mathbb{E}_{\mathcal{S}\sim\Pi(\cdot\mid{\mathbf{y}},i)}\mathbb{E}_{v\sim p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}}),{\mathbf{z}}={\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}}}\left[\mathrm{BCE}(\,\mathbf{1}[v={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{y}})}})\right]$
$=$ $\mathbb{E}_{y,i}\mathbb{E}_{\mathcal{S}\sim\Pi(\cdot\mid{\mathbf{y}},i)}\left[-{\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}q}\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})})-({\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}1-q})\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}1-g_{\phi}^{i}({\mathbf{y}})})\right],$
[…]
$\mathcal{L}(\phi)$ $=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\left[{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})}}\,)}\right]=\mathbb{E}_{y,i}\left[-{\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}q^{\prime}}\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi}^{i}({\mathbf{y}})})-({\color[rgb]{0.3828125,0.2421875,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.3828125,0.2421875,0.6015625}1-q^{\prime}})\log({\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}1-g_{\phi}^{i}({\mathbf{y}})})\right],$
[…]
Discussion on a new notion of per-token quality. Proposition 2 guarantees that the minimizer of the extended PRISM loss satisfies
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi^{\star}}^{i}({\mathbf{y}})}=\mathbb{E}_{\mathcal{S}\sim\Pi(\cdot\mid{\mathbf{y}},i)}\left[p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\mid{\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}})\right].$
Note that by construction, we always have $i\in\mathcal{S}$. This encompasses the $k=1$ case in Section 3.1; When $k=1$, the posterior distribution $\Pi(\cdot\,|\,{\mathbf{y}},i)$ collapses to a point mass at $\mathcal{S}=\{i\}$, recovering Proposition 1.
However, we emphasize that this extended definition still captures our intended notion of per-token quality—namely, the (posterior) likelihood of ${\mathbf{y}}^{i}$ given the rest of the context, here defined under a certain distribution over ${\mathbf{y}}\oplus\mathbf{m}_{\mathcal{S}}$. (note that always $i\in\mathcal{S}$). Consequently, a model fine-tuned with the (extended) PRISM loss retains the capacity to identify low-quality tokens, while allowing $k>1$ to average over multiple plausible masked-context views of $\mathbf{y}$}.
$f_{\theta}$’s effect on PRISM. We show in Proposition 1 and Proposition 2 that, as intended, the design of (b), in which we use $f_{\theta}$ to obtain $\mathbf{y}$}, does not directly influence our theoretical results. To clarify, when $k=1$, the minimizer $g_{\phi^{\star}}^{i}({\mathbf{y}})\;=\;p\!\left({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,\middle|\,{\mathbf{y}}\oplus\mathbf{m}_{i}\right)$ remains fully independent from $f_{\theta}$ (or the rule for choosing $i$), and when $k>1$, the only change is indirect: $f_{\theta}$ (and the rule for choosing index set $\mathcal{S}$ during PRISM fine-tuning) affects the posterior over masked sets $\Pi(\cdot\mid{\mathbf{y}},i)$.
Practically, $f_{\theta}$ shapes the training distribution of samples $({\mathbf{y}},i)$ observed during PRISM fine-tuning, potentially aligning it with those encountered at inference. Since the PRISM loss supervises ${\mathbf{y}}^{i}$, which is sampled from $f_{\theta}$, it is likely to be seen at the inference time (since we also use $f_{\theta}$ during inference to sample clean tokens). We hypothesize that this alignment improves PRISM fine-tuning, and we revisit this point with an ablation on OpenWebText (Appendix D.4).
Appendix C Employing PRISM at inference
In this section, we provide details and pseudocode of inference algorithms that we use for our experiments. We first recall from Section 3.3 how PRISM fine-tuned model ($f_{\theta},g_{\theta}$) is used at inference: At each step $t_{\ell}$, we obtain ${\mathbf{x}}_{t_{\ell+1}}$ through the following steps.
  • (a)
    Choose a subset of masked tokens $\mathcal{S}$ to unmask.
  • (b)
    Choose a subset of clean tokens $\mathcal{T}$ to remask with the lowest quality scores $g_{\theta}({\mathbf{x}}_{t_{\ell}})$.
  • (c)
    For each $i\in\mathcal{T}$, remask ${\mathbf{x}}_{t_{\ell}}^{i}$ and for each $j\in\mathcal{S}$, unmask ${\mathbf{x}}_{t_{\ell}}^{j}$ to a clean token sampled from $f_{\theta}^{j}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})\in\Delta(\mathcal{V})$.
We now provide the pseudocode for two different cases of PRISM implementation, depending on how we choose the sizes of the unmasking and remasking sets $|\mathcal{S}|$ and $|\mathcal{T}|$. We begin with the case where we stochastically assign $|\mathcal{S}|$ and $|\mathcal{T}|$ from binomial distributions (Algorithm  2), corresponding to our experiments on OpenWebText (Section 4.2).
Algorithm 2 PRISM inference with assignments from binomial distributions
1:Require: Fine-tuned MDM with unmasking posterior model $f_{\theta}$, per-token quality head $g_{\theta}$, discretized time steps $\{t_{\ell}\}_{\ell=0}^{N}$, sampling steps NN, sequence length $L$, remasking token ratio $\eta>0$.
2:Initialize ${\mathbf{x}}_{1}={\mathbf{x}}_{t_{1}}\leftarrow(\mathbf{m},\dots,\mathbf{m})$
3:for $\ell=0$ to N1N-1
4:  $K\sim\mathrm{Binom}(|\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}\}|\,,\,\eta)$ and define $\mathcal{M}\leftarrow\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}=\mathbf{m}\}$
5:  if $K\leq|\mathcal{M}|<L-K$ and $l\geq l_{on}$:
6:   $|\mathcal{T}|\leftarrow K$, $|\mathcal{S}|\leftarrow\lceil\frac{L}{N_{s}}\rceil+K$
7:  else:
8:   $|\mathcal{T}|\leftarrow 0$, $|\mathcal{S}|\leftarrow\lceil\frac{L}{N_{s}}\rceil$
9:  Sample $\mathcal{S}\subset\mathcal{M}$ according to the unmasking rule
10:  Sample $\mathcal{T}\subset\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}\}$ with low-$|\mathcal{T}|$ indices of $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{\cdot}(x_{t_{\ell}})$}
11:  Unmask ${\mathbf{x}}^{i}_{t_{\ell}}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})}$ for each $i\in\mathcal{S}$
12:  Remask ${\mathbf{x}}^{j}_{t_{\ell}}$ for each $j\in\mathcal{T}$
13:  ${\mathbf{x}}_{t_{\ell+1}}\leftarrow{\mathbf{x}}_{t_{\ell}}$
14:return ${\mathbf{x}}_{t_{N}}={\mathbf{x}}_{0}$
Now, we state the algorithm where we pre-allocate $|\mathcal{S}|$ and $|\mathcal{T}|$ before inference (Algorithm 3) with fixed $K$, corresponding to our experiments on Sudoku 4.1 and LLaDA 4.3.
Algorithm 3 PRISM inference with fixed $K$
1:Require: Fine-tuned MDM with unmasking posterior model $f_{\theta}$, per-token quality head $g_{\theta}$, sampling steps NN, sequence length $L$, discretized time steps $\{t_{\ell}\}_{\ell=0}^{N}$, number of remasking tokens $K$
2:Initialize ${\mathbf{x}}_{1}={\mathbf{x}}_{t_{1}}\leftarrow(\mathbf{m},\dots,\mathbf{m})$
3:for $l=0$ to N1N-1
4:  Define $\mathcal{M}\leftarrow\{i\mid{\mathbf{x}}_{t_{l}}=\mathbf{m}\}$
5:  if $K\leq|\mathcal{M}|<L-K$ and $l\geq l_{on}$:
6:   $|\mathcal{T}|=K$, $|\mathcal{S}|=\lceil\frac{L}{N}\rceil+K$
7:  else:
8:   $|\mathcal{T}|=0$, $|\mathcal{S}|=\lceil\frac{L}{N}\rceil$
9:  Sample $\mathcal{S}\subset\mathcal{M}$ according to the unmasking rule
10:  Sample $\mathcal{T}\subset\{i\mid{\mathbf{x}}_{t_{\ell}}^{i}\neq\mathbf{m}\}$ with low-$|\mathcal{T}|$ indices of $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{\cdot}(x_{t_{\ell}})$}
11:  Unmask ${\mathbf{x}}^{i}_{t_{\ell}}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{t_{\ell}})}$ for each $i\in\mathcal{S}$
12:  Remask ${\mathbf{x}}^{j}_{t_{\ell}}$ for each $j\in\mathcal{T}$
13:  ${\mathbf{x}}_{t_{\ell+1}}\leftarrow{\mathbf{x}}_{t_{\ell}}$
14:return ${\mathbf{x}}_{t_{N}}={\mathbf{x}}_{0}$
C.1 PRISM-Loop
As we mentioned in Section 4.2, the sentence diversity is crucial for unconditional generation, we observe that the vanilla PRISM inference from Section 3.3 can degrade diversity. we therefore employ PRISM-loop, analogous to ReMDM-loop (wang2025remasking), where we extend the standard generation process with a terminal refinement phase. After the initial generation of a sequence, an additional $N_{l}$ looping steps are performed. In each of these steps, a fixed-size subset of $M$ tokens is randomly selected from the full sequence, remasked, and subsequently unmasked to allow for iterative refinement.
Algorithm 4 PRISM inference with loop strategy
1:Require: Fine-tuned MDM with unmasking posterior model $f_{\theta}$, per-token quality head $g_{\theta}$, sampling steps NN, sequence length $L$, discretized time steps $\{t_{\ell}\}_{\ell=0}^{N}$, remasking token ratio $\eta>0$, loop steps $N_{l}$, number of remasking tokens at the loop stage $M$
2:Sample ${\mathbf{x}}_{0}$ from ${\mathbf{x}}_{1}$ for NN steps with PRISM with $\eta$
3:for $l=1$ to $N_{l}$
4:  Uniformly sample size $M$ subset $\mathcal{T}\subset\{0,\dots,L-1\}$
5:  Remask ${\mathbf{x}}^{i}_{0}$ for each $i\in\mathcal{T}$
6:  Unmask ${\mathbf{x}}^{i}_{0}$ to $v^{i}\sim{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{x}}_{0})}$ for each $i\in\mathcal{T}$
7:return ${\mathbf{x}}_{0}$
Appendix D Experiment details and ablation studies
This section provides additional experimental details.
D.1 Omitted Results
Here, we present the comprehensive results on OpenWebText, on MAUVE score, Generative Perplexity (Gen PPL), and Entropy. The graph in Figure 4 is derived from Table 2, PRISM-loop for $N>128$ and PRISM for $N=128$.
Table 2: Evaluation result for PRISM-loop/PRISM/baselines. The best MAUVE scores are bolded.
Method MAUVE ($\uparrow$) Gen PPL. ($\downarrow$) Entropy ($\uparrow$)
Data 1.00 14.8 5.44
N=1024 N=2048 N=4096 N=1024 N=2048 N=4096 N=1024 N=2048 N=4096
MDLM 0.042 0.037 0.035 51.3 51.3 50.9 5.46 5.46 5.45
ReMDM-conf 0.051 0.060 0.065 48.0 44.5 42.0 5.44 5.42 5.39
ReMDM 0.403 0.610 0.656 28.6 22.8 17.6 5.38 5.30 5.20
[…]
N=128 N=256 N=512 N=128 N=256 N=512 N=128 N=256 N=512
MDLM 0.015 0.023 0.031 61.5 55.8 53.0 5.52 5.49 5.48
ReMDM-conf 0.015 0.022 0.037 74.2 66.0 52.5 5.57 5.54 5.49
ReMDM 0.057 0.216 0.350 42.5 30.5 21.1 5.43 5.34 5.21
[…]
N=16 N=32 N=64 N=16 N=32 N=64 N=16 N=32 N=64
MDLM 0.002 0.006 0.011 230.2 100.5 72.1 5.71 5.60 5.55
ReMDM-conf 0.005 0.006 0.009 244.4 128.4 90.3 5.73 5.66 5.60
ReMDM 0.004 0.005 0.011 257.2 126.4 81.3 5.74 5.65 5.58
[…]
D.2 Details on the remasking Strategy for OpenwebText
This section provides a detailed elaboration of the sampling strategies employed for the unconditional text generation in Section 4.2. The main motivation for these interventions is to preserve sentence diversity, which is crucial for evaluated metrics.
  • Staged Remasking: We restrict the self-correction mechanism to be activated only after a specific step $\ell_{on}$. This restriction is motivated by the observation that low per-token quality scores in the initial sampling phase ($t>t_{\ell_{on}}$) are likely due to a lack of suffix context rather than token-level inaccuracies. Activating the correction mechanism only after a sufficient context prevents spurious modifications to an incomplete sequence.
  • $\mathbf{\eta}$ Inverse Scheduling: Fixing $\eta$ was found to degrade the diversity as the $N$ increases. To mitigate this, an inverse scheduling rule was adopted, dynamically adjusting $\eta$ based on $N$.
To avoid complicated hyperparameter designs, we set these hyperparameters to be a function of $N$. The remasking ratio and the remasking activation step are set to $\eta=\frac{2.56}{N}$ and $l_{on}=\lceil 0.5N\rceil$, respectively. The $N$-step PRISM-loop algorithm comprises an initial $N/2$-step generation for the PRISM and then followed by an $N/2$-step loop with $M=20$.
D.3 Ablation on Fine-tuning Hyperparameters-$(k,n_{y})$
Figure 5: Ablation study on fine-tuning hyperparameters $k$ and $n_{y}$ while holding their product constant ($k\times n_{y}=32$). Evaluations are conducted in ($128,256,512,1024$) sampling steps.
In addition to $k$ (the number of tokens updated simultaneously to sample $\mathbf{y}$} from $\mathbf{z}$}), we introduce a second hyperparameter, $n_{y}\in\mathbb{N}$, to improve data efficiency during the adapter’s fine-tuning. As discussed in Section 3.2, we can generate multiple distinct target sequences $\mathbf{y}$} from a single source sequence $\mathbf{z}$} by using different unmasking index sets $\mathcal{S}$. Therefore, $n_{y}$ denotes the number of unique unmasking sets applied per $\mathbf{z}$} in a single forward pass.
We conducted an ablation study to analyze how the interplay between $k$ and $n_{y}$ affects the generation performance. We tested variants of $(k,n_{y})\in\{(4,8),(8,4),(16,2),(32,1)\}$, thus remaining the total number of token updates per batch constant ($k\times n_{y}=32$) for a fair comparison. All other configurations follow Table  3, except that we set nucleus $p=0.9$ to sample $\mathbf{y}$} from $\mathbf{z}$} (during the PRISM fine-tuning).
Our primary finding is that training the adapter with a smaller number of updating tokens (low $k$) results in a model that performs better at inference time, as measured by the MAUVE score. This can be attributed to a train-test distribution mismatch induced by large $k$. Specifically, updating many tokens simultaneously is likely to introduce joint dependency errors, increasing the chance of sampling a misleading sequence $\mathbf{y}$}. This forces the adapter to train on a corrupted data distribution that is rarely encountered during inference time. Consequently, this mismatch leads to a poorly calibrated quality estimator, degrading the model’s ability to perform effective self-correction.
D.4 Ablation on Fine-tuning Hyperparameters-Nucleus sampling
Figure 6: Ablation study on nucleus sampling $p$ used to generate target sequences during adapter fine-tuning. Evaluations are conducted in ($128,256,512,1024$) sampling steps.
We further investigate how the sampling distribution used to generate target sequences $\mathbf{y}$} affects the adapter’s fine-tuning process. Specifically, we analyze the impact of nucleus sampling probability $p$ that we use to sample a token ${\mathbf{y}}^{i}$ from $f_{\theta}^{i}(\cdot\,|\,{\mathbf{z}})$. Our hypothesis is that a less restrictive sampling distribution (large $p$) is more beneficial for training the quality estimator than a shrunk one (small $p$).
Intuitively, a large $p$ induces a relatively diverse sample $\mathbf{y}$}. These sequences include not only high-probability tokens but also rare yet plausible alternatives. Training on these plausible-but-imperfect examples reduces exposure bias by forcing the model to calibrate its confidence across a wider range of outcomes, rather than only learning from its own single best predictions. This process cultivates a more robust quality estimator capable of discerning subtle inaccuracies. While this approach may introduce gradients with larger variance, it provides superior coverage of the complex distributions encountered during inference. This ultimately leads to a more effective self-correction mechanism and improved text quality, as validated by the results in Fig. 6.
D.5 Experiment Configurations
In Table 3, we list the hyperparameter configurations for Sudoku, OWT, and LLaDA experiments.
Table 3: Hyperparameter configurations for Sudoku, OWT, and LLaDA.
Hyperparameter Sudoku OWT LLaDA
Base Model
   Architecture DiT DiT (non-causual) Transfomer
   Parameters 28.6M 169M 8B
   Tokenizer N/A GPT-2 LLaDA-8B-Instruct
   Sequence Length 81 1024 4096
Fine-tuning
   Adapter Architecture linear attention+linear projection+linear
   Adapter Parameters 133K 7.9M 240M
   Optimizer AdamW AdamW AdamW
   Learning Rate $3.0\times 10^{-4}$ $1.5\times 10^{-4}$ $1.0\times 10^{-4}$
   Weight Decay 0.0 0.0 0.1
   Gloabl Batch Size 256 256 144
   $k$ 4 8 8
   $n_{y}$ 1 4 1
   Nucleus pp 1.0 1.0 0.0
   $\lambda$ 5.0 0.5 0.5
Sampling
   unmasking strategy random random semi-autoregressive
   $n_{\text{blocks}}$ 1 1 32
   Nucleus pp 1.0 0.9 0.0
   Sampling Precision float64 float64 float32
   $K$ 4 N/A
   $l_{on}$ 0 $\lceil 0.5N\rceil$
   $\eta$ N/A $\frac{2.56}{N}$ N/A
D.6 Additional experimental details on LLaDA
Although Table 3 lists the basic configurations for our LLaDA experiments, we provide comprehensive details below.
Attaching the adapter. As discussed in Section 4.3, we freeze the LLaDA-8B-Instruct backbone and add (i) an auxiliary head to model per-token quality and (ii) a LoRA adapter on the qkv attention projections (rank $256$, dropout ratio $0.1$). This yields roughly 250250M trainable parameters in total.
Dataset construction. For PRISM fine-tuning, we adopt the opc-sft-stage-2 (Huang2024OpenCoderTO), comprising $\approx 0.1$M challenging Python coding problems. Each example includes instructions and solutions. Thus, in the masking process, we do not mask instruction tokens.
Training configuration. We use AdamW (loshchilov2017decoupled) (learning rate $1.0\times 10^{-4}$, weight decay $0.01$, warmup ratio $0.05$) and use a cosine schedule with $5$ cycles. Training for $100$ epochs takes approximately $30$ hours on $12$ H100 GPUs.
Inference details. For unmasking, we follow nie2025large and employ semi-autoregressive inference: the sequence is partitioned into multiple blocks, and decoding proceeds from the leftmost block. We set the nucleus parameter to $p=0.0$. These two choices are known to be important for competitive coding performance.
For re-masking, across all baselines, we re-mask exactly $K_{\text{block}}$ tokens per block and schedule these re-masking operations at intermediate stages of block-wise inference (we observe degraded performance if re-masking is concentrated only at the beginning or end of a block). For a fair comparison, we report the best performance for PRISM and baselines (ReMDM, ReMDM-conf) over a sweep of $K_{\text{block}}$ at a fixed number of inference steps $N$: specifically, $K_{\text{block}}\in\{4,6\}$ for $N=256$, (this is because the number of unmasking steps equals $8$ for $N=256$) and $K_{\text{block}}\in\{8,10,12,14,16\}$ for $N=512$ and $N=1024$.